# Simplified and Generalized Masked Diffusion for Discrete Data

### v2

@@ Abstract
Masked (or absorbing) diffusion is actively explored as an alternative to autoregressive models for generative modeling of discrete data. However, existing work in this area has been hindered by unnecessarily complex model formulations and unclear relationships between different perspectives, leading to suboptimal parameterization, training objectives, and ad hoc adjustments to counteract these issues. In this work, we aim to provide a simple and general framework that unlocks the full potential of masked diffusion models. We show that the continuous-time variational objective of masked diffusion models is a simple weighted integral of cross-entropy losses. Our framework also enables training generalized masked diffusion models with state-dependent masking schedules. When evaluated by perplexity, our models trained on OpenWebText surpass prior diffusion language models at GPT-2 scale and demonstrate superior performance on 4 out of 5 zero-shot language modeling tasks. Furthermore, our models vastly outperform previous discrete diffusion models on pixel-level image modeling, achieving 2.75 (CIFAR-10) and 3.40 (ImageNet 64$\times$64) bits per dimension that are better than autoregressive models of similar sizes.
@@ 2 Masked Diffusion
Assume our data lives in a finite discrete state space of size $m$. We use the integers $10,1,\dots,m-$ 1 and their corresponding one-hot vectors $e_{0},e_{1},\dots,e_{m-1}$ to represent the states. [^1] We make no assumption on the ordering of states — any permutation of the integers will give the same result. In masked diffusion, we augment the space with an additional mask state, which is assigned the index $m$. A forward “masking” process is defined such that a data point is evolved into the mask state at some random time. Reversing this process then gives a generative model of the discrete data.
@@
The forward process is defined as a Markovian sequence of discrete random variables $x_{t}$ indexed by time $t$, where $t$ runs from 0 to 1. Throughout the work, we abuse the notation such that $x_{t}$ can be either an integer or its corresponding one-hot vector, whenever it is clear from the context. We divide $[0,1]$ into $T$ intervals, and let $s(i)=(i-1)/T$, $t(i)=i/T$. Following Austin et al. [14], the state transition between $[s(i),t(i)]$ is determined by a transition matrix of size $(m+1)\times(m+1)$: $Q_{i}=(1-\beta_{i})I+\beta_{i}\mathbf{1}e_{m}^{\top},$ where $\mathbf{1}$ is an all-one vector of size $m+1$, $e_{m}$ represents a one-hot vector where element at index $m$ is 1. Each entry $[Q_{i}]_{jk}$ denotes the probability of transition from the state $j$ to the state $k$:
@@ 3 Model and Objective
A single step of our training algorithm is described in Alg. 1 in Appendix.
@@ 4 Sampling
Figure 2: Left: FID evaluation for 50K samples randomly generated from our model on pixel-level modeling of ImageNet 64$\times$64. Right: Number of tokens revealed in each step of generation.
@@ 5 Relation to Existing Work
where $Q(t)$ and $R^{x_{0}}(t)$ are known as the transition rate matrices. Austin et al. [14] derived the same $Q(t)$ in App. A.6 of their paper. However, they did not explore the reverse process or a continuous-time objective. Campbell et al. [29] established an alternative expression for the ELBO using the rate matrices. In Sec. H.1, we show how to recover their expression by separating out a constant from our ELBO expression (4) and applying a discrete “integration-by-part”. A key limitation of their expression is that it needs $N$ evaluations of the prediction model $\mu_{\theta}(\cdot,t)$ to compute an inner summation. To circumvent this computational burden, they used a doubly stochastic estimate. However, this leads to significantly higher variance compared to the analytic cross-entropy (4) which only requires one pass of $\mu_{\theta}(\cdot,t)$. Please refer to Sec. H.2 for more details.
@@
The continuous-time reverse process of our masked diffusion model can be viewed as an any-order autoregressive model (AO-ARM) [46]. To see this, we reorder the tokens according to the timing of their unmasking events in the reverse process. For all tokens, the cumulative distribution functions (CDFs) of unmasking times $\{\tau_{n}\}_{n=1}^{N}$ are identical and satisfy $P(\tau_{n}\leq t)=P(x_{t}^{(n)}=m)=1-\alpha_{t}$. As a result, the ordering is uniformly random across all possible arrangements, and the token prediction during each unmasking event represents a prediction step in AO-ARMs. This connection was initially pointed out in Hoogeboom et al. [47, App. C]. The relation between our simplified ELBO (5) and the AO-ARM objective is independently clarified by Ou et al. [36]. Despite this equivalence, our work demonstrates that the masking schedule $\alpha_{t}$ introduces a new degree of freedom in the design of such models. Variations in $\alpha_{t}$ can lead to different distributions of unmasking times, significantly impacting performance in diffusion-style parallel sampling under time discretization, as shown in Fig. 2.
@@ 6 Generalization to State-dependent Masking Schedules
We consider here state-dependent masking schedules by making the time-dependent probability of masking a token dependent also on the token value. Thus, some tokens will have higher probability to be masked earlier in time compared to others.
@@
This suggests that the reverse model given $x_{t}=m$ can be chosen as $p_{\theta}(x_{s}|x_{t}=m)\triangleq q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mu_{\theta}(x_{t},t)},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{diag}(\mu_{\theta}(x_{t},t))})$ where $\mu_{\theta}(x_{t},t)$ is a probability vector defined as in Sec. 3 that approximates $\mathbb{E}[x_{0}|x_{t}]$ while $\mathrm{diag}(\mu_{\theta}(x_{t},t))$ approximates $\mathbb{E}[x_{0}x_{0}^{\top}|x_{t}]=\mathrm{diag}(\mathbb{E}[x_{0}|x_{t}])$. We show in Sec. I.1 that the negative continuous-time ELBO for the state-dependent rate case is
$\mathcal{L}_{\infty}=\int_{0}^{1}\Big{(}\frac{\alpha_{t}^{\prime}}{\mathbf{1}-\alpha_{t}}\Big{)}^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\left[\delta_{x_{t},m}\cdot(x_{0}-\mu_{\theta}(x_{t},t)+x_{0}x_{0}^{\top}\log\mu_{\theta}(x_{t},t))\right]\mathrm{d}t,$ (12)
where $\alpha_{t}^{\prime}$ is the elementwise derivative of $\alpha_{t}$ w.r.t. $t$. This generalizes the loss from (4), which can be recovered by choosing $\alpha_{t}$ as a scalar schedule times an all-one vector. For $N$ tokens the reverse model and the loss further generalize similarly to Sec. 3; see Sec. I.2.
@@ 7.1 Text
Text is natural discrete data with rich structures. We experiment with two datasets: text8 [54], a character-level text modeling benchmark extracted from English Wikipedia, and OpenWebText [55], an open clone of the unreleased WebText dataset used to train GPT-2 [56]. We choose the linear masking schedule described in Sec. 2 for all text experiments since we find it works best. We refer to our simple masked diffusion model as MD4 (Masked Discrete Diffusion for Discrete Data) and our generalized state-dependent model as GenMD4.
@@
We train a GPT2-small and GPT2-medium size of MD4 on OpenWebText (98% training, 2% validation) and evaluate zero-shot perplexity on the test splits of five benchmark datasets used in Radford et al. [56]: LAMBADA [57], WikiText2, Penn Treebank [58], WikiText103 [59], and One Billion Words [60]. We keep our evaluation setup the same as SEDD [32]. To ensure the comparison is fair, we reproduced their result in our implementation.
The results are shown in LABEL:tab:owt-zeroshot-ppl. Our small model outperforms previous best discrete diffusion models on all five tasks. We also confirmed that our re-implementation of SEDD leads to similar and slightly better results than those reported in their paper. We believe this is because they were using mixed-precision training (thus larger variances) while we are using full precision (float32) training. We are also better than GPT-2 on all tasks except LAMBADA where we are the second best among all methods. When we scale up the model size to medium, we observe that MD4 similarly beats SEDD and GPT-2 on all tasks except LAMBADA.
To validate that the strong zero-shot performance is a result of a better-trained model, we plot perplexity on the OpenWebText validation set in Fig. 4. It demonstrates that our model converges faster and has better final likelihoods than prior methods. We also observed that SEDD [32] has training instabilities. We believe this is due to the score parameterization breaking consistency between forward and reverse processes, as explained in Sec. 5. Although GenMD4 achieves lower perplexity than MD4, we observed that the learned $w$s can easily overfit to dataset statistics, making it less effective on zero-shot transfer tasks. Beyond likelihood evaluation, we also examine generation quality of our models. Fig. 3 showcases a randomly selected, notably coherent sample produced by MD4-medium, alongside its iterative denoising process. Fig. 9 demonstrates MD4’s ability to perform text infilling and highlights a substantial improvement in sample quality when transitioning from the linear to the cosine schedule.
We follow prior work to employ the GPT-2 Large model to evaluate perplexity of the generated samples. As noted in Sec. 4, the generative perplexity metric can be misleadingly decreased by reducing sample diversity; we include these results in Appendix (Fig. 7) mainly for completeness. Under this metric, MD4 significantly outperforms the GPT-2 checkpoint and SEDD. Notably, after 64 decoding steps, our small-sized MD4 matches the performance of SEDD-medium.
@@
Following the convention in prior work [14, 32], we experimented with training masked diffusion models on text chunks of length 256. We used the same dataset splits and transformers of the same size to parameterize the prediction model. Details can be found in Sec. J.1. Results are summarized in Tab. 3. We used the standard bits-per-character metric (a normalized negative log likelihood) for this dataset. We can see that our models outperform previous diffusion models, whether discrete or continuous in nature. We also outperform the best any-order autoregressive models, which do not assume a fixed generation order and has a strong connection to discrete diffusion [47]. Our model is only beaten by an autoregressive transformer and the Discrete Flow (which also used an autoregressive backbone). We believe this is because autoregressive models only require learning a fixed order of generation thus better utilize model capacity.
The text8 dataset has a small vocabulary that consists of 26 letters a-z and a space token. Therefore, we did not expect much flexibility introduced by our state-dependent formulation. Still, with the generalized objective in (12), GenMD4 is able to achieve significantly better BPC than MD4, showing the promise of a state-dependent diffusion on discrete data tasks.
Table 2: Bits Per Character (BPC) on Text8 test set. All models use standard 12-layer transformers similar to GPT2-small [56] except Discrete Flow which uses $8\times 3$ layers.
@@ 7.2 Pixel-level image modeling
To demonstrate MD4 beyond text domains, we follow Austin et al. [14], Campbell et al. [29] and train MD4 on order-agnostic image data. Specifically, we take the CIFAR-10 and the Downsampled ImageNet 64$\times$64 [64] datasets and treat each image as a set of discrete tokens from a vocabulary of size $256$ such that the model is unaware of the relative proximity between different pixel values. We compare to other discrete diffusion and autoregressive models that have reported likelihood results on the same dataset, although to our knowledge there are no published result on discrete diffusion methods for ImageNet $64\times 64$ that directly model raw pixel space.
Results are summarized in Tab. 3. We establish a new state-of-the-art for discrete diffusion models, outperforming previous discrete-time [14] and continuous-time [29] models by a significant margin. Notably, despite lacking knowledge of the ordinal structure of pixel values, MD4 still outperforms models trained with this inductive bias, including D3PM Gauss and $\tau$LDR [29] where the noising distribution is a discrete Gaussian distribution that gives larger probabilities to near pixel values. To isolate the differences caused by training objectives, we also implemented the objective from Campbell et al. [29] in our code. The result, labeled as $\tau$LDR Mask, demonstrates that the high variance of this objective negatively impacts model learning, even when using the same architecture as ours. Our CIFAR-10 model without data augmentation also achieves better likelihood than the previous best autoregressive models on this task, while on ImageNet $64\times 64$ our result is competitive with Transformer AR models which is approximately 4×4$\times$4 × larger in model size and the continuous diffusion model VDM [33].
We provide a random sample from our ImageNet 64$\times$64 model in Fig. 5. More results can be found in in App. K. In Fig. 2, we plot the FID values of samples generated under different choices of schedules and discretization grids. We can see that the model with the linear schedule and a cosine grid achieves an FID close to the model with cosine schedule, both significantly outperform the linear schedule with a uniform grid. We further trained a class-conditional model on ImageNet 64$\times$64 that boosts the FID to around 7. Although these are not state-of-the-art FIDs on ImageNet 64$\times$64, we emphasize our models are optimized for likelihood instead of sample quality.

### v4

@@ Abstract
Masked (or absorbing) diffusion is actively explored as an alternative to autoregressive models for generative modeling of discrete data. However, existing work in this area has been hindered by unnecessarily complex model formulations and unclear relationships between different perspectives, leading to suboptimal parameterization, training objectives, and ad hoc adjustments to counteract these issues. In this work, we aim to provide a simple and general framework that unlocks the full potential of masked diffusion models. We show that the continuous-time variational objective of masked diffusion models is a simple weighted integral of cross-entropy losses. Our framework also enables training generalized masked diffusion models with state-dependent masking schedules. When evaluated by perplexity, our models trained on OpenWebText surpass prior diffusion language models at GPT-2 scale and demonstrate superior performance on 4 out of 5 zero-shot language modeling tasks. Furthermore, our models vastly outperform previous discrete diffusion models on pixel-level image modeling, achieving 2.75 (CIFAR-10) and 3.40 (ImageNet 64$\times$64) bits per dimension that are better than autoregressive models of similar sizes. Our code is available at https://github.com/google-deepmind/md4.
@@ 2 Masked Diffusion
Consider a sentence where we progressively replace each word with a special mask token, transforming the sentence into a sequence of masks. Our goal is to train a generative model that reverses this process, effectively turning a sentence of masks back into meaningful text. More formally, assume our data consists of tokens from a finite discrete state space with $m$ possible states, represented by integers $10,1,\dots,m-$ 1 and their corresponding one-hot vectors $e_{0},e_{1},\dots,e_{m-1}$. To accommodate the masking process, we augment this space with an additional mask state, denoted by the index $m$. The masking process transitions each token to the mask state at a random time. This process, known as the forward process, is applied independently to each token (e.g., each word), progressively converting the data into a sequence of mask tokens. By learning to reverse this masking process, we create a generative model capable of producing coherent discrete data.
@@
We start with the case of a single token and later expand to multiple dimensions. We define the forward process as a Markovian sequence of discrete random variables $x_{t}$ indexed by time $t$, where $t$ runs from 0 to 1. Throughout the work, we abuse the notation such that $x_{t}$ can be either an integer or its corresponding one-hot vector, whenever it is clear from the context. We divide $[0,1]$ into $T$ intervals, and let $s(i)=(i-1)/T$, $t(i)=i/T$. Following Austin et al. [14], the state transition between $[s(i),t(i)]$ is determined by a transition matrix of size $(m+1)\times(m+1)$: $Q_{i}=(1-\beta_{i})I+\beta_{i}\mathbf{1}e_{m}^{\top},$ where $\mathbf{1}$ is an all-one vector of size $m+1$, $e_{m}$ represents a one-hot vector where element at index $m$ is 1. Each entry $[Q_{i}]_{jk}$ denotes the probability of transition from the state $j$ to the state $k$:
@@ 3 Model and Objective
We term our simple masked diffusion model trained with loss (5) MD4 (Masked Discrete Diffusion for Discrete Data). A single step of MD4 training algorithm is described in Alg. 1 in Appendix.
@@ 4 Sampling
Figure 2: Left: FID evaluation for 50k samples randomly generated from MD4 on pixel-level modeling of ImageNet 64$\times$64 (numbers in Tab. 6). Right: Number of tokens revealed per generation step ($T=256$). Each image consists of $64\times 64\times 3=12288$ tokens.
@@ 5 Relation to Existing Work
where $Q(t)$ and $R^{x_{0}}(t)$ are known as the transition rate matrices. Austin et al. [14] derived the same $Q(t)$ in App. A.6 of their paper. However, they did not explore the reverse process or a continuous-time objective. Campbell et al. [29] derived an alternative ELBO expression using rate matrices, which Kitouni et al. [46] further simplified for absorbing diffusion. In Sec. H.1, we show how to recover their expression by separating out a constant from our ELBO expression (4) and applying a discrete “integration-by-part”. A key limitation of their expression is that it needs $N$ evaluations of the prediction model $\mu_{\theta}(\cdot,t)$ to compute an inner summation. To circumvent this computational burden, they used a doubly stochastic estimate. However, this leads to significantly higher variance compared to the analytic cross-entropy (4) which only requires one pass of $\mu_{\theta}(\cdot,t)$. Please refer to Sec. H.2 for more details.
@@
The continuous-time reverse process of our masked diffusion model can be viewed as an any-order autoregressive model (AO-ARM) [47]. To see this, we reorder the tokens according to the timing of their unmasking events in the reverse process. For all tokens, the cumulative distribution functions (CDFs) of unmasking times $\{\tau_{n}\}_{n=1}^{N}$ are identical and satisfy $P(\tau_{n}\leq t)=P(x_{t}^{(n)}=m)=1-\alpha_{t}$. As a result, the ordering is uniformly random across all possible arrangements, and the token prediction during each unmasking event represents a prediction step in AO-ARMs. This connection was initially pointed out in Hoogeboom et al. [48, App. C]. The relation between our simplified ELBO (5) and the AO-ARM objective is independently clarified by Ou et al. [36]. Despite this equivalence, our work demonstrates that the masking schedule $\alpha_{t}$ introduces a new degree of freedom in the design of such models. Variations in $\alpha_{t}$ can lead to different distributions of unmasking times, significantly impacting performance in diffusion-style parallel sampling under time discretization, as shown in Fig. 2.
@@ 6 Generalization to State-dependent Masking Schedules
Consider a scenario where some tokens hold more significance than others and we would like to unmask them earlier in the process. To achieve this, we introduce state-dependent masking schedules, where the probability of unmasking a token depends not only on time, but also on the token’s value.
@@
This suggests that the reverse model given $x_{t}=m$ can be chosen as $p_{\theta}(x_{s}|x_{t}=m)\triangleq q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mu_{\theta}(x_{t},t)},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{diag}(\mu_{\theta}(x_{t},t))})$ where $\mu_{\theta}(x_{t},t)$ is a neural network that approximates $\mathbb{E}[x_{0}|x_{t}]$ while $\mathrm{diag}(\mu_{\theta}(x_{t},t))$ approximates $\mathbb{E}[x_{0}x_{0}^{\top}|x_{t}]=\mathrm{diag}(\mathbb{E}[x_{0}|x_{t}])$. We show in Sec. I.1 that the negative continuous-time ELBO for the state-dependent rate case is
$\mathcal{L}_{\infty}=\int_{0}^{1}\Big{(}\frac{\alpha_{t}^{\prime}}{\mathbf{1}-\alpha_{t}}\Big{)}^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\left[\delta_{x_{t},m}\cdot(x_{0}-\mu_{\theta}(x_{t},t)+x_{0}x_{0}^{\top}\log\mu_{\theta}(x_{t},t))\right]\mathrm{d}t.$ (12)
Here, $\alpha_{t}^{\prime}$ is the elementwise derivative of $\alpha_{t}$. This generalizes the MD4 loss (4), which is recovered when $\alpha_{t}$ is a scalar schedule times a vector of ones. For $N$ tokens, the model further generalize similarly to Sec. 3 and the loss is given in (32). We call this generalized model GenMD4.
@@ 7.1 Text
Text is natural discrete data with rich structures. For comparison with prior work, we evaluate likelihood on two datasets: text8 [55], a character-level text modeling benchmark, and OpenWebText [56], an open clone of the unreleased WebText dataset used to train GPT-2 [57]. We also assess our model’s performance on downstream tasks by training on FineWeb-Edu [58], a high-quality dataset of fine educational text commonly used by the open-source community for comparing LLMs. Unless otherwise specified, a linear schedule and a cosine sampling grid are employed.
@@
We train MD4 of GPT-2 small (S) and GPT-2 medium (M) sizes on OpenWebText and evaluate zero-shot perplexity on five benchmark datasets used in Radford et al. [57]. We keep our evaluation setup the same as SEDD [32]. To ensure fair comparison, we reimplemented SEDD in our codebase. Our implementation led to slightly better results than those reported in their paper.
As seen in LABEL:tab:owt-zeroshot-ppl, our small model outperforms previous best discrete diffusion models on all five tasks. We are also better than GPT-2 on all tasks except LAMBADA where we are the second best method. When scaling up to medium size, MD4 similarly beats SEDD and GPT-2 on 4 out of 5 tasks.
To confirm that the strong zero-shot performance stems from improved training, we plot perplexity on 2% OpenWebText validation set in Fig. 4. Our models converge faster and have better final likelihoods than prior methods. We also observed that SEDD [32] has training instabilities, likely due to score parameterization breaking consistency between forward and reverse processes (Sec. 5). Although GenMD4 achieves lower perplexity than MD4, we observed that the learned $w$s can overfit to dataset statistics, making it less effective on zero-shot transfer tasks.
We also assess our models’ generation quality. Fig. 3 shows a randomly selected, notably coherent sample from MD4-medium and its denoising process. Fig. 10 demonstrates MD4’s text infilling ability and highlights a substantial quality gain when transitioning from uniform to cosine discretization (see Sec. 4). Despite MD4’s strong performance on quantitative metrics like generative perplexity, we have placed these results in Appendix Fig. 8 due to the metric’s inherent unreliability, as noted in Sec. 4. We emphasize the more reliable FID-based assessments found in our image experiments.
@@
Following prior work [14, 32], we trained masked diffusion models on text8 and evaluate the bits-per-character on the test set (details in Sec. J.1). As seen in Tab. 3, our models outperform previous discrete and continuous diffusion models, as well as state-of-the-art AO-ARMs which are closely related to discrete diffusion [48]. Our model is only beaten by an autoregressive (AR) transformer and the AR-backbone Discrete Flow [59]. We believe this is because AR models only require learning a fixed generation order thus better utilize model capacity. Text8’s small vocabulary (26 letters and a space) led us to expect limited flexibility from our state-dependent formulation. However, using the generalized objective in (12), GenMD4 achieved significantly better BPC than MD4, demonstrating the potential of state-dependent diffusion for discrete data.
Figure 5: Hellaswag accuracy vs. training steps for MD4 and AR models at GPT-2 small, medium, and large scales.
FineWeb-Edu.
We train MD4 on FineWeb-Edu and evaluate its zero-shot accuracy on the Hellaswag dataset [60], a popular common sense inference benchmark for LLMs. We directly compared MD4 to its AR counterparts – transformers with identical configurations (except for causal masking) trained on the same data. Results are summarized in Fig. 5.
MD4 demonstrates steady performance growth with increasing scale. While outperformed by AR models of the same size, the performance gap does not widen as model size increases. For example, AR-small reaches 30% accuracy in 50k steps, while MD4-small takes 200k steps (4x data efficiency difference). At the medium scale, AR achieves 37% in 270k steps, compared to MD4’s 1 million steps.
Table 2: Bits Per Character (BPC) on Text8 test set. All models use standard 12-layer transformers similar to GPT-2 small [57] except Discrete Flow which uses $8\times 3$ layers.
@@ 7.2 Pixel-level image modeling
Unlike continuous diffusion which struggles with discrete data, we show that MD4, a discrete diffusion model, performs well on inherently continuous data, suggesting its potential for unifying modalities. We follow Austin et al. [14] and train MD4 on order-agnostic image data from CIFAR-10 and downsampled ImageNet 64$\times$64 [63]. Each image is treated as a set of 256-valued discrete tokens, making the model agnostic to pixel proximity. We compare to other discrete diffusion and AR models with reported likelihood results on these datasets, although to our knowledge there are no published result on discrete diffusion for ImageNet $64\times 64$ that directly model raw pixel space.
Tab. 3 summarizes our results. We establish a new state-of-the-art for discrete diffusion models, outperforming previous work [14, 29] by a significant margin. Our CIFAR-10 result surpasses the best reported AR result. On ImageNet $64\times 64$, our results are competitive with Transformer AR models that are 4$\times$ larger, as well as a strong continuous diffusion model VDM [33]. Notably, despite lacking knowledge of the ordinal structure of pixel values, MD4 outperforms models trained with this inductive bias, including D3PM Gauss and Campbell et al. [29] where the noising distribution is a discrete Gaussian that assigns larger probabilities to near pixel values. To isolate the differences caused by training objectives, we also implemented the Campbell et al. [29] objective with the absorbing process, showing its high variance hinders learning even with our architecture.
We provide a random sample from our ImageNet 64$\times$64 model in Fig. 6. More results can be found in App. K. In Fig. 2, we plot the FID values of samples generated under different choices of schedules and discretization grids. We can see that the model with the linear schedule plus a cosine grid achieves an FID close to the model with cosine schedule, both significantly outperform the linear schedule with a uniform grid. We further trained a class-conditional model on ImageNet 64$\times$64 that boosts the FID to around 7. Although these are not state-of-the-art FIDs on ImageNet 64$\times$64, we emphasize our models are optimized for likelihood instead of sample quality.